OpenStudy (anonymous):

amount of wax needed to make 200 cylindricle candles, given :- length of candle=12CM, base radius = 2cm

OpenStudy (diyadiya):

Volume of 1 candle = pi*r^2*h =pi*4*12=48pi wax for 200 candles=200*48*pi =9600pi =3.14*9600 =30144 cm^3

hero (hero):

Diyadiya, I saw you

OpenStudy (diyadiya):

Lol i was just checking if i went wrong or not :D

hero (hero):

I hadn't bothered to attempt this.

OpenStudy (diyadiya):

lol ok

hero (hero):

It looks right

OpenStudy (diyadiya):

okay :)

hero (hero):

Diyadiya, what is your relation to Lana?

OpenStudy (diyadiya):

hm what ?

hero (hero):

nvm

hero (hero):

efc?